OpenStudy (anonymous):

what is the domain and range of (absolute value of sinx)^2

OpenStudy (solomonzelman):

\(\large\color{black}{ \left| \sin x \right| ^2 }\). Well, for all we know, \(\large\color{black}{ -1\le \sin x\le1 }\). (this is also true about the cosine of x) So, the range of \(\large\color{black}{ \left| \sin x \right| \le1 }\) would be just \(\large\color{black}{ [0,1] }\) . \(\normalsize\color{black}{ \bullet }\) Because negatives are not allowed, so \(\normalsize\color{blue}{ \rm from~zero }\). \(\normalsize\color{black}{ \bullet }\) \(\normalsize\color{blue}{ \rm Till~1 }\), Because sin(x) can't be equal to a number greater than 1.

OpenStudy (solomonzelman):

the range of \(\large\color{black}{ \left| \sin x \right|^2 }\) would not differ (from \(\large\color{black}{ \left| \sin x \right| }\) ). (the \(\large\color{black}{ \left| \sin x \right|^2 }\) would have corners but \(\large\color{black}{ \left| \sin x \right| }\) would not)

OpenStudy (solomonzelman):

Now, the domain can be any number, because there is no restriction on what you are plugging in for x. \(\large\color{black}{ \left| \sin x \right|^2 }\) is defined for all values of \(\large\color{black}{ x }\). So the domain is \(\large\color{black}{ (-\infty ,+\infty) }\).